In an Earth reference frame, a star is 76 light-years away. How fast would you have to travel so that to you the distance would be only 39 light-years?
Answer #1

sqrt(1-beta^2) = L1 / L2 = 39/76 = 0.513

(1-beta^2)= 0.2634

beta^2= 0.736

beta = sqrt(0.736) =0.858

v = 0.858 C where C = 3e8 m/s